You want to put $8,000 in a simple interest account. It has a 2% annual interest rate. How long must you invest that money to earn $400 in interest?

Found 2 solutions by Theo, Edwin McCravy:
Answer by Theo(13342)

the simple interest formula is i = p * r * n
p is the principal
r is the interest rate per time period
n is the number of time periods.

in your problem, p is 8000 and r is .02 and i is 400.
your time period is years.
formula becomes 400 = 8000 * .02 * n
simplify to get 400 = 160 * n
solve for n to get n = 400 / 160 = 2.5 years.
8000 * .02 * 2.5 = 400
that's your solution.
the money must be invested for 2.5 years.

Answer by Edwin McCravy(20055)


  I = (P)(R)(T)

400 = (8000)(0.02)(T) 

400 = 160(T)

400   160T
——— = ————
160   160

2.5 = T

Two and a half years.

Edwin